What is the maximum amount you would pay for an asset that generates an income of $250,000 at the end of each of five years, if the opportunity cost of using funds is 8 percent?

Instructions: Do not round intermediate calculations. Round only your final calculation to the nearest penny (two decimal places). Use a negative sign (-) where appropriate.

$ _____________
Answer #1

Answer

The maximum price is equal to the present value of the cash flows

PV=A*(P|A,i,n)

A=annual cashflow

i=8%

n=5 years

PV=250000*(P|A,8%,5)

=250000*3.992710037078

=998177.509

the maximum price you will pay is $998177.51